In the gaseous reactiojn 2A+B = A2B, Gibbs free energy=1200cal at 227Degree Celsius. What total pressure would be necessary to produce 60% conversion of B into A2B when 2:1 mixture is used?
Ans (5.02atm)
Sir, it is the question of "Modern Approach To chemical calculations" by Ramendra c. Mukerjee, page no.507, Q.no.31.

Grade:Upto college level

1 Answers

Askiitians expert Anirudh iitk
14 years ago

delG = -RT ln K.

Find out K

initially:

say x moles of B is present, then 2x moels of A is present. Let total pressure be P.

Finally:

say a moles of B reacted. Then. 2a mol of A reacts and a mol of A2B is formed.

 

given : a=0.6x.

Hence, B- 0.4x and A - 1.2x and A2B= 0.6x

partial pressure of A2B is (0.6/(0.6+1.2+0.4))*P = (0.6/2.2)P

similarly P= (0.4/2.2)P, Pa = (1.2/2.2)P

 

Kp = Pa2b / Pb .Pa 2    This value of Kp is obtained using the topmost formula. From this find out the value of P and that is the required value